What is the empirical formula if 38.23 % Oxygen, 50.61 % Cu, and 11.16 % Nitrogen? [empForm]


1
Expert's answer
2021-02-24T05:31:20-0500

Assume 100 g of the substance.

Convert the grams of each element to moles by dividing by their molar mass:

n(O) = 38.23 g / 16 g/mol = 2.4 mol

n(Cu) = 50.61 g / 64 g/mol = 0.8 mol

n(N) = 11.16 g / 14 g/mol = 0.8 mol

Divide each of the three mole figures by the lowest of the three in order to simplify the mole ratio:

n(O) = 2.4 mol / 0.8 mol = 3

n(Cu) = 0.8 mol / 0.8 mol = 1

n(N) = 0.8 mol / 0.8 mol = 1

As a result, the empirical formula is CuNO3.


Answer: CuNO3